OpenStudy (anonymous):

an airplane travels 1,800 miles in 3 hours flying with the wind. On the return trip, flying against the wind it takes 4 hours to travel 2,000 miles. Find the rate of the plane in still air.

OpenStudy (anonymous):

The rate of the plane in still air is 550 mph.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Use Distance = Rate * Time 1800 = (x + c) * 3 2000 = (x - c) * 4

OpenStudy (anonymous):

x = the rate in still air c = the rate of the wind

OpenStudy (anonymous):

1800 = 3x + 3c 2000 = 4x - 4c

OpenStudy (anonymous):

1800 - 3x = 3c 600 - x = c

OpenStudy (anonymous):

2000 = 4x - 4 (600 - x) 2000 = 4x - 2400 + 4x 2000 = 8x -2400 4400 = 8x 550 = x

OpenStudy (anonymous):

thanks